A long segment of text can also be emphasized in this way. Text within such a segment given additional emphasis with Roman type. Italic type loses its ability to emphasize and become simply distracting when used excessively.

It is sometimes necessary to prevent TEX from breaking a line where it might otherwise do so. This may be at a space, as between the “Mr.” and “Jones” in “Mr. Jones”, or within a word—especially when the word is a symbol like itemnum that makes little sense when hyphenated across lines.

Footnotes^{1} pose no problem.

TEX is good at typesetting mathematical formulas like x − 3y = 7 or a_{1} > x^{2n}/y^{2n} > x′. Remember that a letter like x is a formula when it denotes a mathematical symbol, and should be treated as one.

2 Displayed Text

Text is displayed by indenting it from the left margin. Quotations are commonly displayed. There are short quotations

This is a short a quotation. It consists of a single paragraph of text.

There is no paragraph indentation.

and longer ones.

This is a longer quotation. It consists of two paragraphs of text. The beginning of each paragraph is indicated by an extra indentation.

This is the second paragraph of the quotation. It is just as dull as the first paragraph.

Another frequently-displayed structure is a list. The following is an example of an itemized list.

• This is the first item of an itemized list. Each item in the list is marked with a “tick”. The document style determines what kind of tick mark is used.

• This is the second item of the list. It contains another list nested inside it. The inner list is an enumerated list.

1. This is the first item of an enumerated list that is nested within the itemized list.

2. This is the second item of the inner list. LaTEX allows you to nest lists deeper than you really should.

This is the rest of the second item of the outer list. It is no more interesting than any other part of the item.

1This is an example of a footnote.

• This is the third item of the list.

You can even display poetry.

There is an environment for verse Whose features some poets will curse.

For instead of making Them do all line breaking, It allows them to put too many words on a line when they’d rather

be forced to be terse.

Mathematical formulas may also be displayed. A displayed formula is oneline long; multiline formulas require special formatting instructions.

x′ + y^{2} = z_{i}^{2}

Don’t start a paragraph with a displayed equation, nor make one a paragraph by itself.